OpenStudy (flygirlmandi):

convert the map scale to a unit rate. how many inches represent one mile? interpret the meaning of the unit rate. 3/10 inches = 7/8 miles thanks :)

563blackghost (563blackghost):

3/10 in = 7/8 mi multiply both sides by 10 3/10 * 10 in = 7/8 * 10 mi 3 in = 70/8 mi divide both sides by 3 1 in = 70/8 / 3 = 70/24 = 35/12 mi ~cited from http://openstudy.com/study#/updates/528182f7e4b00cd889de86f1

OpenStudy (flygirlmandi):

thank you :)

563blackghost (563blackghost):

np ^^

563blackghost (563blackghost):

plz close your question once you have attained your answer ^^